# Predicting the outcome of collisional neutrino flavor conversion

### v2

@@ III.3 Prediction of the asymptotic state
Our goal is to analytically predict the values taken by $N$ and $\overline{N}$ for $t\to\infty$, which we will denote with a superscript ()($\infty$)( ∞ ). More specifically, in the spirit of designing a semi-classical subgrid model which provides a prescription for the evolution of the classical neutrino distributions (so without treating flavor coherence, even though it is crucially nonzero in the final state), we want to predict the asymptotic values of the flavor-diagonal components $N_{aa}$ and $\overline{N}_{aa}$. As indicated by the numerical results of Fig. 2, this asymptotic state is characterized by: (i) constant flavor-diagonal number densities, (ii) a configuration lying at the edge of collisional instability. Interestingly, a similar “edge of instability” configuration was obtained in [75] for a system externally driven to a (fast) instability, and is reminiscent of more general considerations on the equilibrium of a neutrino system, as seen in the framework of plasmon-like collective flavor conversions (see, e.g., [76, 77, 78]) or miscidynamics [79, 65].
As stated above, the vacuum term contribution is negligible for the non-linear evolution, such that $N_{xx}\simeq\overline{N}_{xx}$ at all times. Steady-state equations for condition (i) can be written as $\mathrm{Tr}(\dot{N})=0$ and $\dot{N}_{ee}-\dot{\overline{N}}_{ee}=0$, that is,
@@
Condition (ii) reads differently depending on which mode was unstable. If the system was in a “minus” instability, the final state satisfies $\mathrm{Im}(\Omega_{-}^{(\infty)})=0$, that is,
@@ III.4 Magnitude of flavor coherence
In the previous section, we have focused on the quantum steady-state reached by the flavor on-diagonal components of $N$. We now provide a more complete picture of this asymptotic state, discussing flavor coherence. To that end, we introduce the modulus and phase of the flavor off-diagonal component of the density matrix, specifically:
@@
$\left[R_{\pm}^{(\infty)}\right]^{2}\simeq\frac{2\,\alpha_{\pm}\,\kappa_{x}}{(\kappa_{e}-\bar{\kappa}_{e})^{2}}\,\mathrm{Im}(\Omega_{\pm})\,\left[N_{ee}^{(\mathrm{cl})}-\overline{N}_{ee}^{(\mathrm{cl})}\right]^{2}\,,$ (18)
with $\alpha_{+}=1$ and $\alpha_{-}=\kappa_{e}/\bar{\kappa}_{e}$, and the growth rate of the instability is given by Eq. (7). Since by definition, $\mathrm{Im}(\Omega_{\pm})>0$ if there is an instability, this ensures that the right-hand side of Eq. (18) is positive.
@@ III.5 Duration of relaxation
but stopped at $2N_{xx}^{(\infty)}-N_{xx}^{(=)}$ (symmetric value from $N_{xx}^{(=)}$ with respect to $N_{xx}^{(\infty)}$). The duration of this phase is thus [^3] We approximate this portion of exponential relaxation at the rate $\kappa_{x}$ by its linear counterpart, i.e., $e^{-\kappa_{x}\Delta t}\simeq 1-\kappa_{x}\Delta t$.
@@
which gives $\Delta t\simeq 180\,\mathrm{\mu s}$ for both points A and B, a value consistent with Fig. 2. One should note, however, that these are long timescales on which the matter field could evolve, such that our effectively homogeneous treatment (see discussion at the end of Sec. II.1) could be challenged.
@@ IV Importance of classical relaxation
For Point B, the evolution in the $\kappa_{x}=0$ case is similar to that of Point A: after the saturation of the instability, $N_{xx}$ remains stuck on $N_{xx}^{(=)}=\overline{N}_{ee}^{(\mathrm{cl})}$, and $N_{ee}$ and $\overline{N}_{ee}$ are back to their classical equilibrium values. The evolution in the “off-diagonal damping only” case is more subtle. At first, following the plus instability, there is a flavor swap between $\nu_{x}$ and $\bar{\nu}_{e}$ (see the almost perfect inversion of the dash-dotted blue and green lines) in the first $100\,\mathrm{\mu s}$. However, this new configuration is unstable, but for the minus mode, while the plus mode is stable (with a negative imaginary part). This is the reason why $|N_{ex}|$ first decreases, as the plus mode decays, before the newly unstable minus mode takes over and leads to a second instability, this time inducing the equipartition of $\nu_{x}$ and $\nu_{e}$. With the “full” collision term, this change of unstable mode is prevented by the fact that $N_{ee}$ and $\overline{N}_{ee}$ remain close to their classical equilibrium values. Interestingly, the outcome (swap/equipartition) of the “off-diagonal damping only” CFI is related to which mode (plus/minus) is unstable, as noted in [66]—although that work used multi-energy distributions.
@@ Acknowledgements.
I thank J. Kneller, G. McLaughlin, S. Richers, F. Foucart, E. Grohs, C. Hall and E. Urquilla for many useful discussions, and in particular J. Kneller, G. McLaughlin, S. Richers, H. Nagakura and D. Fiorillo for their valuable comments on the manuscript. This work was supported by the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center award No. PHY-2020275.
@@ Appendix A Linear stability analysis
Throughout this section, we drop the ${(\mathrm{cl})}$ exponents which are implicit.
Neglecting the vacuum contribution, the stability matrix in Eq. (4) can be rewritten
$\Lambda=\begin{bmatrix}A-G-\mathrm{i}\,(\alpha+\gamma)-\omega\,c_{2\theta}+\lambda&A+G\\ A-G&A+G-\mathrm{i}\,(\alpha+\gamma)+\omega\,c_{2\theta}+\lambda\end{bmatrix}\,.$ (23)
@@
$Q_{1}$ $=\begin{bmatrix}G+\mathrm{i}\alpha+\sqrt{A^{2}+2\mathrm{i}G\alpha-\alpha^{2}}\\ G-A\end{bmatrix}\,,$ $\ \Omega_{1}$ $=\lambda+A-\sqrt{A^{2}+(\mathrm{i}\alpha+\omega c_{2\theta})(2G+\mathrm{i}\alpha+\omega c_{2\theta})}-\mathrm{i}\gamma\,,$ (24)
$Q_{2}$ $=\begin{bmatrix}G+\mathrm{i}\alpha-\sqrt{A^{2}+2\mathrm{i}G\alpha-\alpha^{2}}\\ G-A\end{bmatrix}\,,$ $\ \Omega_{2}$ $=\lambda+A+\sqrt{A^{2}+(\mathrm{i}\alpha+\omega c_{2\theta})(2G+\mathrm{i}\alpha+\omega c_{2\theta})}-\mathrm{i}\gamma\,.$
Outside of the resonance-like regime, we have $\lvert G\alpha\rvert\ll A^{2}$ (and also $|\alpha|\ll G$, $|\omega c_{2\theta}|\ll G$), which allows to simplify the expressions as
@@ Appendix B Attenuation of the Hamiltonian
As mentioned in Sec. III.2, for computational reasons we use an attenuated Hamiltonian, where the vacuum, matter and self-interaction contributions in Eq. (1) are multiplied by an attenuation factor [^6] This approach is slightly different from the one in, e.g., [28, 32], where only the self-interaction Hamiltonian $H_{\nu\nu}$ is rescaled. $\eta<1$. This is necessary in practice because the matter term in particular introduces an extremely small timescale (see Table 2). In addition, the matter term reduces the amplitude of the eigenvectors $Q_{\pm}$ in Eq. (6), see Eq. (12) in [54] with the expressions of $\Omega_{\pm}$ given in Eqs. (27) and (29), thus requiring exceedingly small error tolerances in the ODE solver.
@@
We observe from Fig. 5 that using $\eta>10^{-4}$ does not change the evolution and in particular the final state equilibrium; we are then safely in the regime where the classical relaxation timescale is much larger than the self-interaction (and matter) timescales. Ideally, one could circumvent this procedure by tracking only the “slow” evolution driven by collisions (which have secular effects on the very fast oscillations happening on matter/self-interaction timescales), following strategies which leverage the large timescale separation, such as the adiabatic-like approximation in [81, 82], the coarse-graining of [65] or the “slow-dynamics equations” derived in [64].
Figure 5: Modification of the evolution for different attenuation factors $\eta$ in the Hamiltonian (see text). There is no discernible difference for attenuations larger than $10^{-4}$, where we are safely in the regime $\Gamma\ll||H||$.

### v3

@@ III.3 Prediction of the asymptotic state
Our goal is to analytically predict the values taken by $N$ and $\overline{N}$ for $t\to\infty$, which we will denote with a superscript ()($\infty$)( ∞ ). More specifically, in the spirit of designing a semiclassical subgrid model which provides a prescription for the evolution of the classical neutrino distributions (so without treating flavor coherence, even though it is crucially nonzero in the final state), we want to predict the asymptotic values of the flavor-diagonal components $N_{aa}$ and $\overline{N}_{aa}$. As indicated by the numerical results of Fig. 2, this asymptotic state is characterized by (i) constant flavor-diagonal number densities, and (ii) a configuration lying at the edge of collisional instability. Interestingly, a similar “edge of instability” configuration was obtained in [75] for a system externally driven to a (fast) instability, and is reminiscent of more general considerations on the equilibrium of a neutrino system, as seen in the framework of plasmonlike collective flavor conversions (see, e.g., [76, 77, 78]) or miscidynamics [79, 65].
As stated above, the vacuum term contribution is negligible for the nonlinear evolution, such that $N_{xx}\simeq\overline{N}_{xx}$ at all times. Steady-state equations for condition (i) can be written as $\mathrm{Tr}(\dot{N})=0$ and $\dot{N}_{ee}-\dot{\overline{N}}_{ee}=0$, that is
@@
Condition (ii) reads differently depending on which mode was unstable. If the system was in a “minus” instability, the final state satisfies $\mathrm{Im}(\Omega_{-}^{(\infty)})=0$, that is
@@
Generalizations —
Recent simulations (see e.g., [80, 81, 82, 83, 84]) have shown that muons could be present and have a significant impact on the dynamics in CCSNe or NSMs. For such situations, we could not assume $N_{xx}\simeq\overline{N}_{xx}$ and would have to consider a full 4-species treatment. The derivation of the asymptotic state is straightforward to generalize, with $[\overline{N}_{ee}^{(\infty)}-N_{xx}^{(\infty)}]\to[\overline{N}_{ee}^{(\infty)}-\overline{N}_{xx}^{(\infty)}]$ in Eqs. (9)–(10) and with the additional (now independent) equation $\mathrm{Tr}(\dot{\overline{N}})=0$, reading $\bar{\kappa}_{e}[\overline{N}_{ee}^{(\infty)}-\overline{N}_{ee}^{(\mathrm{cl})}]=-\bar{\kappa}_{x}[\overline{N}_{xx}^{(\infty)}-\overline{N}_{xx}^{(\mathrm{cl})}]$ (note that now $\bar{\kappa}_{x}\neq\kappa_{x}$). Although such a configuration is outside the scope of this paper and has not been numerically studied, we postulate that these small changes would describe the CFI asymptotic state with $\nu_{x}\neq\bar{\nu}_{x}$.
Strictly speaking, four species would not be enough and one should consider in general configurations where $\nu_{\mu}\neq\bar{\nu}_{\mu}\neq\nu_{\tau}\neq\bar{\nu}_{\tau}$, which we leave for future work. We note that the two-flavor scheme developed here could still be useful to determine a three-flavor asymptotic state, similarly to the method employed in [85] in the fast regime.
@@ III.4 Magnitude of flavor coherence
In the previous section, we focused on the quantum steady-state reached by the flavor on-diagonal components of $N$. We now provide a more complete picture of this asymptotic state, discussing flavor coherence. To that end, we introduce the modulus and phase of the flavor off-diagonal component of the density matrix, specifically,
@@
$\left[R_{\pm}^{(\infty)}\right]^{2}\simeq\frac{2\,a_{\pm}\,\kappa_{x}}{(\kappa_{e}-\bar{\kappa}_{e})^{2}}\,\mathrm{Im}(\Omega_{\pm})\,\left[N_{ee}^{(\mathrm{cl})}-\overline{N}_{ee}^{(\mathrm{cl})}\right]^{2}\,,$ (18)
with $a_{+}=1$ and $a_{-}=\kappa_{e}/\bar{\kappa}_{e}$, and the growth rate of the instability is given by Eq. (7). Since by definition, $\mathrm{Im}(\Omega_{\pm})>0$ if there is an instability, this ensures that the right-hand side of Eq. (18) is positive.
@@ III.5 Duration of relaxation
but this classical relaxation brings the system back into an unstable configuration (see bottom panel of Fig. 2). Flavor coherence grows and flavor conversion $\{\nu_{e},\bar{\nu}_{e}\}\to\nu_{x}$ occurs. Equation (20) describes the evolution of $N_{xx}$ only up to a point, which we take to be roughly symmetric to $N_{xx}^{(=)}$ with respect to $N_{xx}^{(\infty)}$, that is, [^3] We use Eq. (20) from $N_{xx}^{(=)}=N_{xx}^{(\infty)}+\Delta N$ to $N_{xx}^{(\infty)}-\Delta N$, the latter being therefore equal to $2N_{xx}^{(\infty)}-N_{xx}^{(=)}$. $2N_{xx}^{(\infty)}-N_{xx}^{(=)}$. The duration of this phase is thus [^4] We approximate this portion of exponential relaxation at the rate $\kappa_{x}$ by its linear counterpart, i.e., $e^{-\kappa_{x}\Delta t}\simeq 1-\kappa_{x}\Delta t$.
@@
which gives $\Delta t\simeq 180\,\mathrm{\mu s}$ for both points A and B. This value is consistent with the duration of the first “classical relaxation/instability” cycle in Fig. 2. The overall timescale associated to the evolution toward $N^{(\infty)}$ corresponds to a few of these cycles, and can thus be estimated by a few $\Delta t$. One should note, however, that these are long timescales on which the matter field could evolve, such that our effectively homogeneous treatment (see discussion at the end of Sec. II.1) could be challenged.
@@ IV Importance of classical relaxation
For Point B, the evolution in the $\kappa_{x}=0$ case is similar to that of Point A: after the saturation of the instability, $N_{xx}$ remains stuck on $N_{xx}^{(=)}=\overline{N}_{ee}^{(\mathrm{cl})}$, and $N_{ee}$ and $\overline{N}_{ee}$ are back to their classical equilibrium values. The evolution in the “off-diagonal damping only” case is more subtle. At first, following the plus instability, there is a flavor swap between $\nu_{x}$ and $\bar{\nu}_{e}$ (see the almost perfect inversion of the dash-dotted blue and green lines) in the first $100\,\mathrm{\mu s}$. However, this new configuration is unstable, but for the minus mode, while the plus mode is stable (with a negative imaginary part). This is the reason why $|N_{ex}|$ first decreases, as the plus mode decays, before the newly unstable minus mode takes over and leads to a second instability, this time inducing the equipartition of $\nu_{x}$ and $\nu_{e}$. With the “full” collision term, this change of unstable mode is prevented by the fact that $N_{ee}$ and $\overline{N}_{ee}$ remain close to their classical equilibrium values. Interestingly, the outcome (swap/equipartition) of the “off-diagonal damping only” CFI is related to which mode (plus/minus) is unstable, as noted in [66], although that work used multienergy distributions.
@@ Acknowledgements.
I thank J. Kneller, G. McLaughlin, S. Richers, F. Foucart, E. Grohs, C. Hall and E. Urquilla for many useful discussions, and in particular J. Kneller, G. McLaughlin, S. Richers, H. Nagakura, M. Zaizen and D. Fiorillo for their valuable comments on the manuscript. This work was supported by the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center award No. PHY-2020275.
@@ Appendix A Linear stability analysis
The stability matrix in Eq. (4) can be rewritten
$\Lambda=\begin{bmatrix}A-G-\mathrm{i}\,(\gamma+\alpha)-\omega\,c_{2\theta}+\lambda&A+G\\ A-G&A+G-\mathrm{i}\,(\gamma-\alpha)+\omega\,c_{2\theta}+\lambda\end{bmatrix}\,.$ (23)
@@
$Q_{1}$ $=\begin{bmatrix}G+\mathrm{i}\,\alpha_{\omega}+\sqrt{A^{2}+2\,\mathrm{i}\,G\,\alpha_{\omega}-\alpha_{\omega}^{2}}\\ G-A\end{bmatrix}\,,$ $\ \Omega_{1}$ $=\lambda+A-\sqrt{A^{2}+2\,\mathrm{i}\,G\,\alpha_{\omega}-\alpha_{\omega}^{2}}-\mathrm{i}\,\gamma\,,$ (24)
$Q_{2}$ $=\begin{bmatrix}G+\mathrm{i}\,\alpha_{\omega}-\sqrt{A^{2}+2\,\mathrm{i}\,G\,\alpha_{\omega}-\alpha_{\omega}^{2}}\\ G-A\end{bmatrix}\,,$ $\ \Omega_{2}$ $=\lambda+A+\sqrt{A^{2}+2\,\mathrm{i}\,G\,\alpha_{\omega}-\alpha_{\omega}^{2}}-\mathrm{i}\,\gamma\,,$
where we wrote $\mathrm{i}\alpha_{\omega}\equiv\mathrm{i}\alpha+\omega c_{2\theta}$. Outside of the resonancelike regime, we have $\lvert G\alpha\rvert\ll A^{2}$ (and also $|\alpha|\ll G$, $|\omega c_{2\theta}|\ll G$), which allows one to simplify the expressions as
@@
Figure 5: Modification of the evolution for different attenuation factors $\eta$ in the Hamiltonian (see text). There is no discernible difference for attenuations larger than $10^{-4}$, where we are safely in the regime $\Gamma\ll||H||$. For comparison purposes, the calculations are aligned at the saturation time $t_{\mathrm{sat}}$, which corresponds to the first peak in $|N_{ex}|$.
@@ Appendix B Attenuation of the Hamiltonian
As mentioned in Sec. III.2, for computational reasons we use an attenuated Hamiltonian, where the vacuum, matter and self-interaction contributions in Eq. (1) are multiplied by an attenuation factor [^7] This approach is slightly different from the one in, e.g., [28, 32], where only the self-interaction Hamiltonian $H_{\nu\nu}$ is rescaled. $\eta<1$. This is necessary in practice because the matter term in particular introduces an extremely small timescale (see Table 2). In addition, the matter term reduces the amplitude of the eigenvectors $Q_{\pm}$ in Eq. (6), see Eq. (12) in [54] with the expressions of $\Omega_{\pm}$ given in Eqs. (27) and (29), thus requiring exceedingly small error tolerances in the differential equation solver.
@@
We observe from Fig. 5 that using $\eta>10^{-4}$ does not change the evolution and in particular the final state equilibrium; we are then safely in the regime where the classical relaxation timescale is much larger than the self-interaction (and matter) timescales. Using the notations from Appendix A, attenuating the Hamiltonian by $\eta$ corresponds to rescaling $G\to\eta G$ and $A\to\eta A$. Therefore, as long as $\lvert(\eta G)\alpha\rvert\ll(\eta A)^{2}$, the instability growth rate is given by the approximate Eq. (25) where the factors of $\eta$ cancel out, which explains why the growth rate is unchanged in Fig. 5. We have checked that smaller values $\eta<10^{-6}$ (which break the physical ordering of timescales) lead to a different evolution, with in particular a smaller growth rate.
Ideally, one could circumvent this attenuation procedure by tracking only the “slow” evolution driven by collisions (which have secular effects on the very fast oscillations happening on matter/self-interaction timescales), following strategies which leverage the large timescale separation, such as the adiabatic-like approximation in [87, 88], the coarse graining of [65] or the “slow-dynamics equations” derived in [64].